# Online POMDP Planning with Anytime Deterministic Optimality Guarantees

### v4

@@ 2 Related Work
While the primary focus of this paper is on discrete POMDP planning, it is essential to acknowledge recent advancements in POMDP planning that encompass both discrete and continuous observation spaces. Few notable approaches include POMCPOW (Sunberg and Kochenderfer, 2018), LABECOP (Hoerger and Kurniawati, 2021) and AdaOPS (Wu et al., 2021), which leverage explicit use of observation models. These algorithms employ importance sampling mechanisms to weigh each state sample based on its likelihood value, which is assumed to be known. Although these methods have exhibited promising performance in practical scenarios, they currently lack formal guarantees. To address this gap, (Lim et al., 2020, 2022) introduced a simplified solver aimed at bridging the theoretical gap between the empirical success of these algorithms and the absence of theoretical guarantees for continuous observation spaces. In (Lim et al., 2022), probabilistic guarantees were derived for the simplified solver estimating its proximity to the optimal value function.
@@ 3 Preliminaries
A policy function $a_{t}=\pi_{t}(H_{t})$ determines the action to be taken at time step $t$, based on the history $H_{t}$ and time $t$. In the rest of the paper we write $\pi_{t}\equiv\pi_{t}(H_{t})$ for conciseness. The reward is defined as an expectation over a state-dependent function, $r(b_{t},a_{t})=\mathbb{E}_{x\sim b_{t}}[r_{x}(x,a_{t})]$, and is assumed to be bounded by $-\mathcal{R}_{\max}\leq r_{x}(x,a_{t})\leq\mathcal{R}_{\max}$. The value function for a policy $\pi$ over a finite horizon $T$ is defined as the expected cumulative reward received by executing $\pi$ and can be computed using the Bellman update equation,
@@ 4 Simplified POMDP
¯$\bar{\mathbb{P}}(x_{t+1}\mid x_{t},a_{t})\triangleq$$\begin{cases}\mathbb{P}(x_{t+1}\mid x_{t},a_{t})&,\ x_{t+1}\in\bar{\mathcal{X}}(H_{t+1}^{-})\\ 0&,\ otherwise\end{cases}$ (5)
¯$\bar{\mathbb{P}}(z_{t}\mid x_{t})\triangleq$$\begin{cases}\mathbb{P}(z_{t}\mid x_{t})&,\ z_{t}\in\bar{\mathcal{Z}}(H_{t})\\ 0&,\ otherwise\end{cases}$ (6)
@@ 5.1.2 Optimality Guarantees for Simplified Observation Spaces
The proof is provided in 4. ∎
Notably, using UDB to find the optimal policy does not require a recovery of all the observations in the theoretical belief tree, but only a subset which depends on the definition and complexity of the POMDP. Each action-value is bounded by a lower and upper bound, which can be represented as an interval enclosing the theoretical value. When the bound intervals of two candidate actions do not overlap, one can clearly discern which action is suboptimal, rendering its subtree redundant for further exploration. This distinction sets UDB apart from current state-of-the-art online POMDP algorithms. In those methods, any finite-time stopping condition fails to ensure optimality since the bounds used are either heuristic or probabilistic in nature.
@@ 5.2.2 Optimality Guarantees
The proof is provided in 9.5.
In this lemma, we establish that employing the ’partial’ root-bound is sufficient for ensuring both upper and lower bounds in relation to the optimal value function at the root node. This approach differs from that presented in the previous section (see Lemma 1). There, each node in the tree was associated with its unique upper bound based on its value function. In contrast, the current lemma demonstrates that using the ’partial’ bound across all nodes in the tree, which is valid only at the root, still guarantees bounded value for the optimal root-value function, while avoiding the requirement to maintain a complete belief at each node of the tree.
@@ 5.2.3 Early Stopping Criteria
Lemma 2 establishes that the recursive Bellman-like optimality operator, can be used to bound the optimal value function at the root. Since the bounds are deterministic, these bounds can be used for eliminating suboptimal actions with full certainty while planning.
@@ 5.2.4 Exploration Strategies
where the actions are chosen by the highest upper bound, sometimes referred to as an "optimism in face of uncertainty", which offers a balance between exploration and exploitation of actions that are possibly optimal or have high uncertainty in their value. Observations are chosen based on the maximum gap between the upper and lower bounds, which results in observations with high uncertainty in their value. Last, we define ¯$\bar{\mathbb{P}}^{\star}(\tau_{t})$ as the probability of a trajectory $\tau_{t}$ under a policy derived from recursive action selection as per (29). Subsequently, the selection of states effectively maximizes the difference in probability between the individual trajectory density and the aggregate probability of all sampled trajectories that begin with that particular trajectory.
Lemma 3.
Performing exploration based on (29), (30) and (31) ensures that the algorithm converges to the optimal value function within a finite number of planning iterations. The proof is provided in 7.
Importantly, alternative methods for the state-action-observation exploration are viable and, if given limited planning time, may offer improved performance in practice. Lemma 3 suggests one way that is guaranteed to converge in finite time. We leave the investigation of other approaches for finite-time convergence using the deterministic bounds for future research.
@@ 6 Algorithms
6:        ¯$\bar{\mathbb{P}}(h_{0})\xleftarrow{}\bar{\mathbb{P}}(h_{0})+\bar{\mathbb{P}}_{0}$
@@
0:   function fwdUpdate($ha,haz$, $\tau_{d}$, ¯$\bar{\mathbb{P}}_{\tau}$, $x^{\prime}$)
@@
6:  ¯$\bar{\mathbb{P}}_{\tau}\xleftarrow{}\bar{\mathbb{P}}_{\tau}\cdot Z_{z\mid x^{\prime}}\cdot T_{x^{\prime}\mid x,a}$
@@
8:     ¯$\bar{\mathbb{P}}(haz)\xleftarrow{}\bar{\mathbb{P}}(haz)+\bar{\mathbb{P}}_{\tau}$
@@
In this section we aim to describe how to fit our bounds to a blueprint of a general algorithm, named $\textsc{Algorithm}-\mathcal{A}$, which serves as an abstraction to many existing algorithms. Then, we explicitly describe two algorithms, DB-POMCP, an adaptation to POMCP that uses UCB for exploration, and our deterministic bounds for decision-making, and RB-POMCP, a particle-based solver that utilizes the bounds both for decision-making and exploration.
@@
$\textsc{Algorithm}-\mathcal{A}$ is outlined in algorithm 1. For clarity of exposition, we assume the following; at each iteration a single state particle is propagated from the root node to the leaf (line 2 of function Search). The selection of the next state and observations are done by sampling from the observation and transition models (line 5), and each iteration ends with the full horizon of the POMDP (lines 2). However, none of these are a restriction of our approach and may be replaced with arbitrary number of particles, arbitrary state and observation selection mechanism and a single or multiple expansions of new belief nodes at each iteration.
@@ 6.3 Time complexity
The time complexity for each posterior node, primarily depends on the specific algorithm being used. In the case of dynamic programming methods, such as DESPOT and AdaOPS, there is a negligible added computational complexity detailed below. In the case of Monte Carlo methods, such as POMCP, the computational complexity is $O(|\mathcal{A}|)$ attributed mainly to the action-selection, while our approach adds another linear time complexity term, making it $O(|\mathcal{A}|+|\bar{\mathcal{Z}}|)$ due to the summation over the simplified observation space. During each iteration of the algorithm, an "IF" statement is used to determine whether a specific trajectory has already been encountered at the current node. This verification process can potentially result in an added linear complexity of $O(D)$, where $D$ represents the planning horizon. However, this overhead can be circumvented by assigning a unique ID value to each trajectory at the previous step and subsequently checking whether a pair, comprising the ID value and the new state, has already been visited. This approach reduces the overhead to an average time complexity of $O(1)$ by utilizing hash maps efficiently.
@@ 7.1 Deterministic-Bounds for Decision-Making
Our experimental analysis reveals that, in addition to offering a level of optimality certification for the chosen actions, utilizing deterministic bounds for action selection can enhance the expected cumulative reward. It is important to note, however, that this method does not always lead to better outcomes. Specifically, it may not be advantageous in situations where the highest lower bound is less than other available upper bounds (for instance, comparing actions $a^{1}$ and $a^{3}$ in figure 2). This limitation is evident in the results for the Laser Tag POMDP, a considerably larger problem compared to the other POMDPs evaluated, where the deterministic bounds did not yield performance improvements.
@@
Algorithm Tiger POMDP Laser Tag Discrete Light Dark Baby POMDP
[…]
@@ 7.2 Root-Bounds for Decision-Making and Exploration
[…]
The performance outcomes presented in Table 2 reveal that the RB-POMCP algorithm typically matches or surpasses the standard POMCP in various tested environments, except for the Navigate to Goal POMDP scenario. The limited performance in this particular context can be attributed to the nature of RB-POMCP’s exploration strategy, which is designed to assure optimality over extended planning periods but does not inherently guarantee enhanced results within limited planning durations. Unlike probabilistic algorithms that leverage statistical concentration inequalities—such as the Hoeffding inequality employed in the Upper Confidence Bounds for Trees (UCT) Kocsis and Szepesvári (2006) exploration mechanism of POMCP—RB-POMCP adopts a more cautious strategy. This approach entails considering both worst-case and best-case scenarios to establish a deterministic link with the optimal value may not always translate to superior immediate performance due to its conservative nature.

### v5

@@ 2 Related Work
SARSOP, like other point-based offline solvers, provides suboptimality guarantees after completion of an offline solve, whereas our setting is online decision making under tight per-decision budgets, yielding a certificate at each iteration. In large spaces with concentrated posteriors, exhaustive belief updates can allocate computation to low-probability regions; sampling-based online planners (including ours) naturally bias effort toward high-probability trajectories. Consequently, SARSOP is not our primary baseline for time-constrained online planning. That said, its focus on “optimally reachable” beliefs is complementary and can inform our selection of simplified state/observation models without changing guarantees.
@@
While the primary focus of this paper is on discrete POMDP planning, it is essential to acknowledge recent advancements in POMDP planning that encompass both discrete and continuous observation spaces. Few notable approaches include POMCPOW (Sunberg and Kochenderfer, 2018), LABECOP (Hoerger and Kurniawati, 2021) and AdaOPS (Wu et al., 2021), which leverage explicit use of observation models. These algorithms employ importance sampling mechanisms to weigh each state sample based on its likelihood value, which is assumed to be known. Although these methods have exhibited promising performance in practical scenarios, POMCPOW and LABECOP currently lack formal guarantees, while (Wu et al., 2021) derives probabilistic guarantees which do not hold in practice for AdaOPS algorithm, due to assumption relaxations. To address this gap, (Lim et al., 2020, 2023) introduced a simplified solver aimed at bridging the theoretical gap between the empirical success of these algorithms and the absence of theoretical guarantees for continuous observation spaces. (Lim et al., 2023), derive (i) high-probability bounds on the value loss of their particle-belief solver and (ii) a deterministic guarantee on its expected return-obtained after averaging over both the solvers internal randomness and the POMDPs stochastic dynamics. Because this guarantee is only in expectation, a single execution may still violate the bound, whereas the results presented in this paper hold deterministically for every run.
@@ 3 Preliminaries
A policy function $a_{t}=\pi_{t}(H_{t})$ determines the action to be taken at time step $t$, based on the history $H_{t}$ and time $t$. In the rest of the paper we write $\pi_{t}\equiv\pi_{t}(H_{t})$ for conciseness. The reward is defined as an expectation over a state-dependent function, $r(b_{t},a_{t})=\mathbb{E}_{x\sim b_{t}}[r_{x}(x,a_{t})]$, and is assumed to be bounded by $\mathcal{R}_{\max}\leq r_{x}(x,a_{t})\leq\mathcal{R}_{\max}$. The value function for a policy $\pi$ over a finite horizon $T$ is defined as the expected cumulative reward received by executing $\pi$ and can be computed using the Bellman update equation,
@@ 4 Simplified POMDP
$\bar{\mathbb{P}}(x_{t+1}\mid x_{t},a_{t})\triangleq$ $\begin{cases}\mathbb{P}(x_{t+1}\mid x_{t},a_{t})&,\ x_{t+1}\in\bar{\mathcal{X}}(H_{t+1}^{-})\\ 0&,\ otherwise\end{cases}$ (5)
$\bar{\mathbb{P}}(z_{t}\mid x_{t})\triangleq$ $\begin{cases}\mathbb{P}(z_{t}\mid x_{t})&,\ z_{t}\in\bar{\mathcal{Z}}(H_{t})\\ 0&,\ otherwise\end{cases}$ (6)
@@ 5.1.2 Optimality Guarantees for Simplified Observation Spaces
The proof is provided in Section˜A.1.2. ∎
Notably, using UDB to find the optimal policy does not require a recovery of all the observations in the theoretical belief tree, but only a subset which depends on the definition and complexity of the POMDP. Each action-value is bounded by a lower and upper bound, which can be represented as an interval enclosing the theoretical value. When the bound intervals of two candidate actions do not overlap, one can clearly discern which action is suboptimal, rendering its subtree redundant for further exploration. This distinction sets UDB apart from current state-of-the-art online POMDP algorithms. In those methods, any finite-time stopping condition fails to ensure optimality since the bounds used are either heuristic or probabilistic in nature. Note, however, that previous offline algorithms did utilize similar pruning, such as SARSOP.
@@ 5.2.2 Optimality Guarantees
Proof.
The proof is provided in A.1.5. ∎
In this corollary, we establish that employing the ’partial’ root-bound is sufficient for ensuring both upper and lower bounds in relation to the optimal value function at the root node. This approach differs from that presented in the previous section (see Lemma 1). There, each node in the tree was associated with its unique upper bound based on its value function. In contrast, the current corollary demonstrates that using the ’partial’ bound across all nodes in the tree, which is valid only at the root, still guarantees bounded value for the optimal root-value function, while avoiding the requirement to maintain a complete belief at each node of the tree.
@@ 5.2.3 Early Stopping Criteria
Corollary 2.1 establishes that the recursive Bellman-like optimality operator, can be used to bound the optimal value function at the root. Since the bounds are deterministic, these bounds can be used for eliminating suboptimal actions with full certainty while planning.
@@ 5.2.4 Exploration Strategies
where the actions are chosen by the highest upper bound, sometimes referred to as an "optimism in face of uncertainty", which offers a balance between exploration and exploitation of actions that are possibly optimal or have high uncertainty in their value. Observations are chosen based on the maximum gap between the upper and lower bounds, which results in observations with high uncertainty in their value. Last, we define $\bar{\mathbb{P}}^{\star}(\tau_{t})$ as the probability of a trajectory $\tau_{t}$ under a policy derived from recursive action selection as per (29). Subsequently, the selection of states effectively maximizes the difference in probability between the individual trajectory density and the aggregate probability of all sampled trajectories that begin with that particular trajectory.
Corollary 2.2.
Proof.
The proof is provided in A.1.6. ∎
Importantly, alternative methods for the state-action-observation exploration are viable and, if given limited planning time, may offer improved performance in practice. Corollary 2.2 suggests one way that is guaranteed to converge in finite time. We leave the investigation of other approaches for finite-time convergence using the deterministic bounds for future research.
@@
5.3 Impact of POMDP Characteristics on Deterministic Bounds
To make explicit how POMDP characteristics affect bound tightness, we recast the upper–lower gap in terms of simplified-space coverage and local value spread.
The analytical bounds depend on trajectory-probability terms such as $\sum_{\tau}\bar{\mathbb{P}}(\tau)$ and on the auxiliary bounds $\mathcal{V}_{\max,t},\mathcal{V}_{\min,t}$; starting from the recursive definitions in (21), let
$\Delta V(H_{t})\;\triangleq\;\mathcal{V}_{\max}(H_{t})-\mathcal{V}_{\min}(H_{t}),$
$\delta(H_{t})\;\triangleq\;\sum_{\tau_{t}\in\mathcal{T}(H_{t})}\bar{\mathbb{P}}(\tau_{t})\;-\!\!\!\sum_{z_{t+1}\in\bar{\mathcal{Z}}(H_{t},\pi_{t})}\;\sum_{\tau_{t+1}\in\mathcal{T}(H_{t+1})}\bar{\mathbb{P}}(\tau_{t+1}),$
then following the derivation in A.1.7,
$\mathcal{U}^{\pi}_{0}(H_{t})-\mathcal{L}^{\pi}_{0}(H_{t})\;=\;\Delta V(H_{t})\,\delta(H_{t})+\!\!\!\!\!\sum_{z_{t+1}\in\bar{\mathcal{Z}}(H_{t},\pi_{t})}\!\!\!\!\!\left[\mathcal{U}^{\pi}_{0}(H_{t+1})-\mathcal{L}^{\pi}_{0}(H_{t+1})\right].$ (32)
Equation (32) indicates that the gap accumulates the value function spread $\Delta V(H_{k})$ only on the time-steps where the trajectory is not part of the simplified set. The attributes govern the magnitude of the bounds are twofold:
  1. 1.
    Coverage probability of the simplified spaces. An important quantity is the probability mass of trajectories not part of the simplified sets, $\bar{\mathcal{X}}\!,\bar{\mathcal{Z}}$, rather than the cardinality of the omitted sets. A large POMDP whose belief is highly concentrated within the simplified spaces can therefore yield a smaller gap than a small, nearly uniform domain.
  2. 2.
    Magnitude of the local value spread, $\Delta V(H_{k})$. Each time a trajectory leaves the simplified spaces, its contribution to the gap is scaled by $\Delta V(H_{k})=\mathcal{V}_{\max}(H_{k})-\mathcal{V}_{\min}(H_{k})$, the policy-independent spread between the upper and lower cumulative value bounds at node $H_{k}$. Sharpening these auxiliary bounds, e.g. by using informed rollouts or MDP-optimal surrogates, reduces $\Delta V(H_{k})$ and proportionally narrows the overall gap.
Practical implication. Maximizing the probability coverage of the simplified spaces, followed by tightening the auxiliary bounds to reduce $\Delta V(H_{k})$, yields a more favourable gap-complexity trade-off than a uniform expansion of the simplified spaces.
@@ 6 Algorithms
6:    $\bar{\mathbb{P}}(h_{0})\xleftarrow{}\bar{\mathbb{P}}(h_{0})+\bar{\mathbb{P}}_{0}$
@@
0:   function fwdUpdate($ha,haz$, $\tau_{d}$, $\bar{\mathbb{P}}_{\tau}$, $x^{\prime}$)
@@
6:  $\bar{\mathbb{P}}_{\tau}\xleftarrow{}\bar{\mathbb{P}}_{\tau}\cdot Z_{z\mid x^{\prime}}\cdot T_{x^{\prime}\mid x,a}$
@@
8:   $\bar{\mathbb{P}}(haz)\xleftarrow{}\bar{\mathbb{P}}(haz)+\bar{\mathbb{P}}_{\tau}$
@@
[…]
@@
[…]
@@ 6.3 Time complexity
The cost of updating a posterior node depends on the underlying solver. For dynamic programming methods, such as DESPOT and AdaOPS, the extra bookkeeping required by our bounds is negligible, so the original time complexity is essentially unchanged. For Monte Carlo methods, such as POMCP, the baseline complexity is $O(|\mathcal{A}|)$ related mainly to the action-selection; A naive implementation of our bounds adds another linear complexity term, making it $O(|\mathcal{A}|+|\bar{\mathcal{Z}}|)$ due to the summation over the simplified observation space described in Corollary 2.1 and algorithm 1, line 1 of bwdUpdate function.
However, we show that our added bounds can still be updated in $O(|\mathcal{A}|)$ by storing two additional scalar bookkeeping values for each node. Namely, incremental visitation probability, $\Delta\mathbb{P}^{(i)}(\cdot)$, and the change to child’s upper bound, $\Delta\mathcal{U}^{(i)}(\cdot)$.
Each POMCP visit touches a single observation child, so only that branch must be updated. The bounds are therefore updated incrementally,
$\mathcal{U}_{0}^{(i)}(H_{t},a_{t})=\mathcal{U}_{0}^{(i-1)}(H_{t},a_{t})+\Delta\overline{\mathbb{P}}^{(i)}\left(\tau_{t}^{(i)}\right)\left[r\left(x_{t}^{(i)},a_{t}\right)+V_{max,t}\right]$ (34)
$+\Delta\mathcal{U}_{0}^{(i)}\left(H_{t+1}^{(i)}\right)-\Delta\overline{\mathbb{P}}^{(i)}\left(\tau_{t+1}^{(i)}\right)V_{max,t}$
and,
$\mathcal{U}_{0}^{(i)}(H_{T})=\underset{a_{t}\in\mathcal{A}}{\max}\left\{\mathcal{U}_{0}^{(i)}(H_{t},a_{t})\right\}.$ (35)
Because only one branch is modified, the per-visit overhead remains $O(|\mathcal{A}|)$. The full algebraic derivation is provided in Section˜A.1.7.
During each visitation to a node, the trajectory bookkeeping $\tau_{d}\in\tau(ha)$ shown in algorithm 1 line 1, is used to determine whether a specific trajectory has already been encountered at the current node. This verification process can potentially result in an added linear complexity of $O(D)$, where $D$ represents the planning horizon. However, this overhead can be circumvented by assigning a unique ID value to each trajectory at the previous step and subsequently checking whether a pair, comprising the ID value and the new state, has already been visited. This approach reduces the overhead to an average time complexity of $O(1)$ by utilizing hash maps efficiently.
While the asymptotic time complexity remains similar, the deterministic certificates add a constant-factor bookkeeping cost per node, due to hash lookups and updates of bounds ($O(|A|)$ worst-case plus $O(1)$ on average for algorithm 1). Practical wall-clock effects depend on implementation details and engineering optimizations, so we confine this section to the algorithmic analysis; systematic profiling is orthogonal to our contribution and left to future work.
@@ 7.1 Deterministic-Bounds for Decision-Making
Our experimental analysis reveals that, in addition to offering a level of optimality certification for the chosen actions, utilizing deterministic bounds for action selection can enhance the expected cumulative reward. It is important to note, however, that this method does not always lead to better outcomes. Specifically, it may not be advantageous in situations where the highest lower bound is less than other available upper bounds (for instance, comparing actions $a^{1}$ and $a^{3}$ in figure 2). In practice, when the bounds overlap, using them for action selection provides only heuristic guidance and may be incorrect. In our experiments, we used uniform bounds accross all future belief states, which are not useful for guiding the policy towards better actions. In fact, the more symmetric the bounds are at different nodes, the more belief nodes the algorithm will have to visit to distinguish between the attractiveness of different actions. While this can be improved by considering informed bounds, e.g. MDP-optimal for upper bounds or rollout-based lower bounds, we leave it for future work. When using uninformed bounds, a large POMDP with large state, observation or action spaces, leads to more loose bounds, which in turn reduces the effectiveness of our approach. This limitation is evident in the results for the Laser Tag POMDP, a considerably larger problem compared to the other POMDPs evaluated, where the deterministic bounds did not yield performance improvements.
@@
Tiger POMDP Laser Tag Discrete Light Dark Baby POMDP
Cardinalities $(|S|,|A|,|O|)$ (2, 3, 2) (5930, 5, $\sim\!1.5\!\times\!10^{6}$) (122, 5, 21) (2, 2, 2)
DB-DESPOT (ours) $3.74$ p m 0.48 5.29-5.29 p m 0.14 5.29-5.29 p m 0.01 3.92-3.92 p m 0.56
AR-DESPOT $2.82$ p m 0.55 5.10-5.10 p m 0.14 61.53-61.53 p m 5.80 5.40-5.40 p m 0.85
DB-POMCP (ours) $3.01$ p m 0.21 3.97-3.97 p m 0.24 3.70-3.70 p m 0.82 4.48-4.48 p m 0.57
POMCP $2.18$ p m 0.76 3.92-3.92 p m 0.27 4.51-4.51 p m 1.15 5.39-5.39 p m 0.63
@@ 7.2 Root-Bounds for Decision-Making and Exploration
Cardinalities $(|S|,|A|,|O|)$ (2, 3, 2) (1801, 8, 3) (25, 5, 25) (2, 2, 2)
RB-POMCP (ours) $1.53$ p m 0.76 $8.50$ p m 0.22 $61.21$ p m 0.71 11.97-11.97 p m 0.27
DB-POMCP (ours) 1.05-1.05 p m 0.15 $7.86$ p m 0.21 $62.37$ p m 0.75 12.13-12.13 p m 0.22
POMCP 5.59-5.59 p m 0.24 $5.69$ p m 0.20 $68.45$ p m 0.69 12.49-12.49 p m 0.27
The performance outcomes presented in Table 2 reveal that the RB-POMCP algorithm typically matches or surpasses the standard POMCP in various tested environments, except for the Navigate to Goal POMDP scenario. The limited performance in this particular context can be attributed to the nature of RB-POMCP’s exploration strategy, which is designed to assure optimality over extended planning periods but does not inherently guarantee enhanced results within limited planning durations. Unlike probabilistic algorithms that leverage statistical concentration inequalities-such as the Hoeffding inequality employed in the Upper Confidence Bounds for Trees (UCT) Kocsis and Szepesvári (2006) exploration mechanism of POMCP-RB-POMCP adopts a more cautious strategy. This approach entails considering both worst-case and best-case scenarios to establish a deterministic link with the optimal value may not always translate to superior immediate performance due to its conservative nature.